# Influence-aware Causal Autoencoder Network for Node Importance Ranking in Complex Networks

### v1

@@ IV-D Implementation details
The learning rate is $\mu=0.001$ and the maximum number of iterations $T=10$. In addition, we set the recovery probability $\delta=1$ and the infection probability $\gamma=1.5\times\gamma_{c}$ to ensure effective spreading within the network. The infection threshold $\gamma_{c}$ is derived from mean-field theory [51] as
@@
For other comparison methods, all the parameters are set as default. For model evaluation, Kendall’s $\tau$ coefficient on the test networks is adopted as the primary evaluation metric. To reduce the impact of randomness on experimental results, both ICAN and the comparative methods are run 5 times each, with average values recorded. All experiments are conducted on a computer running Windows 10, equipped with an Intel(R) i5-10400 CPU at 2.90 GHz and 16 GB of memory.
IV-E Results on various real-world networks (RQ1)
@@
IV-E1 Training on BA-generated network and testing on various real-world networks
In the first experiment, we conduct experiments by training on the BA generated network and testing on different real-world networks to validate the model’s effectiveness. Table II presents the predicted Kendall’s correlation coefficients across different datasets. The experimental results comprehensively evaluate the proposed ICAN model against several baseline methods for node ranking across six real-world networks. As we can see, ICAN demonstrates superior and consistent performance. Notably, ICAN achieves the highest Kendall’s coefficient on every individual dataset. This culminates in the best overall average Kendall’s coefficient of 0.7707, securing the top rank among all compared methods. The fact that ICAN, trained on a BA synthetic network, generalizes effectively to diverse real-world networks highlights its robust cross-network predictive capability. This can be attributed to ICAN learning network-invariant node embeddings that transfer reliably across different networks.
@@ IV-E1 Training on BA-generated network and testing on various real-world networks
IV-E2 Training on ER-generated network and Testing on various real-world networks
@@ IV-F Ablation study
IV-F1 Ablation study for the influence-aware causal representation learning mechanism (RQ1)
@@
As shown in the equations of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, both are associated with $\bm{W}$, through which the node embeddings are learned to be causally related to the node influence scores. This indicates that the causal reconstruction loss and the causal ranking loss jointly contribute to the process of influence-aware causal representation learning. To emphasize the advantages of this influence-aware causal mechanism in ICAN, we conduct an ablation study by setting $\lambda_{1}=0$ and fixing $\bm{W}$ as an identity matrix. In this setup, the Markov blanket of the node influence score variable in the causal reconstruction loss is replaced by the low-dimensional representation directly obtained from the autoencoder. This variant can be regarded as a classical autoencoder applied to the ranking task and is denoted as ICAN_w/o_Cau. The comparison results between ICAN and ICAN_w/o_Cau across different networks are shown in Fig. 2(a). A noticeable performance decline of ICAN_w/o_Cau is observed in most node ranking tasks, highlighting the critical importance of the causal mechanism integrated into the autoencoder model.
IV-F2 Ablation study for the feature–task co-optimization mechanism (RQ2)
@@
In this study, we integrate the causal ranking loss and causal reconstruction loss within a unified framework to enable causal feature extraction and ranking tasks. To evaluate the effectiveness of the proposed end-to-end feature–task co-optimization mechanism, we conduct an ablation study by transforming the integrated model into a two-stage implementation, in which the feature extraction and ranking processes are decoupled, thereby removing the synergistic optimization effect. This variant is referred to as ICAN_w/o_Syner. In the first stage, the causal reconstruction loss and a regularization term are employed as the loss function to optimize the causal representation learning module, yielding low-dimensional latent causal representations that comprehensively encapsulate the topological properties of the network. During the second stage, the learned causal representations are fed into the ranking prediction module, which is optimized using the causal ranking loss along with a regularization term to generate the predicted node influence score. These predictions are then compared against the node influence score derived from the SIR model, with Kendall’s $\tau$ coefficient used as the evaluation metric. The two-stage model is trained on the BA synthetic network and evaluated on six real-world networks. The comparison results between ICAN and ICAN_w/o_Syner on different networks are presented in Fig. 2(b). Performance decline in ICAN_w/o_Syner is observed in most of the node ranking tasks. These results demonstrate that ICAN facilitates mutual reinforcement between causal representation learning and ranking prediction, resulting in more accurate ranking outcomes.
IV-F3 Ablation study for the causal ranking loss (RQ3)
@@
To validate the efficacy of the proposed the causal ranking loss, we conduct ablation experiments, replacing the CausalListMLE with the ListMLE loss for the node importance ranking , defined as ICAN_w/o_CausalRank, and additionally replacing the CausalListMLE with a MSE loss for the node importance score regression, denoted as ICAN_w/o_Rank. As illustrated in Fig. 2(c), the ICAN model consistently and significantly outperforms the ICAN_w/o_CausalRank and ICAN_w/o_Rank variants on all datasets. The superiority of the CausalListMLE loss stems from its inherent ability to capture ordinal relationships within ranked lists, which is more aligned with the nature of the node importance ranking objective. Overall, these findings conclusively demonstrate that integrating the causal ranking loss loss function is indispensable for achieving high-ranking accuracy, thereby affirming the design rationale behind our task-aware mechanism.
IV-G Degree distribution visualization (RQ4)
In this section, we analyze the degree distribution of nodes identified as important according to the predicted ranking. Figs. 3 and 4 compare the degree distribution of the three networks and the corresponding influence scores of the top 10% nodes identified by ICAN, DC, and CGNN. Each subfigure combines a blue histogram representing the network’s degree distribution (left $y$-axis, relative frequency) with red scatter points indicating the influence scores of individual nodes (right $y$-axis, node influence score).
A clear methodological contrast emerges. As expected for the degree–centrality-based approach, the results for DC in Fig. 3(b) and 4(b) exhibit a strong positive correlation: nodes with the highest influence scores coincide with those of the largest degrees, concentrating red points at the high-degree end of the spectrum. In contrast, ICAN and CGNN reveal markedly different patterns. High-scoring nodes appear across a broader degree range, with a noticeable concentration in the moderate-degree region. This suggests that these methods attribute high influence to nodes that are not necessarily the network’s primary hubs.
This divergence highlights a key insight: ICAN captures influence patterns that go beyond simple degree centrality. By identifying influential nodes through a more nuanced, potentially propagation-aware representation, ICAN emphasizes structurally balanced nodes that may serve as more efficient and cost-effective targets for network interventions.
Figure 3: Degree distribution of Jazz network and top $10\%$ influential nodes.
Figure 4: Degree distribution of USAir network and top $10\%$ influential nodes.
IV-H Experiments with various training networks (RQ5)
@@
IV-I Case study for attribute graphs
Previously, we conducted experiments on six real-world networks without attributes, where node features are generated by Node2Vec. In this case study, we show that ICAN can be applied to an attributed network. The feature matrix can be constructed from node attributes. Cora [52] is a citation network where nodes represent scientific publications and edges represent directed citation links between them. Each node feature in the Cora dataset is a 1,433-dimensional binary bag-of-words vector representing the presence or absence of specific terms in the paper’s content. Our model is trained on BA synthetic networks (using Node2Vec for feature generation) and evaluated on the attributed Cora real-world network. The experimental results, as shown in Fig. 6, demonstrate that our method achieves superior ranking performance compared to all baseline approaches. This indicates that ICAN effectively generalizes across different types of networks and is adept at integrating both topological and attribute information for improved representation learning.
Figure 6: Kendall’s $\tau$ of different methods on Cora dataset.
IV-J Parameter sensitivity analysis
@@
V Related work
Our framework is a node-ranking method built on causal representation learning. Below, we review the two research areas most closely related to our model.
Node ranking methods: Previous works for node importance ranking can be categorized into two main groups: traditional topological-based methods and deep representation learning-based methods. Over the past decades, numerous topology-based node ranking methods have been developed, which can be broadly categorized into neighborhood-based, eigenvector-based, and path-based approaches. Neighborhood-based methods, such as degree centrality [53], H-index [54], and $K$-shell decomposition [49], assess node importance using local structural information. In contrast, eigenvector-based methods like eigenvector centrality [12, 55] and PageRank [56] capture global influence propagation. Path-based methods, including betweenness centrality [46] and eccentricity centrality [57], evaluate nodes based on their positions and roles in network paths. Apart from non-deep techniques, GCN based deep represeantion learning models have shown their superiority over most of existing methods. Zhao et al. [14] proposed the InfGCN algorithm, which takes neighbor graphs and classic structural features as the input into a graph convolutional network for learning nodes’ representations, and then feeds the representations into a classifier layer. Keikha et al. [58] proposed DeepIM that employs network embedding via deep learning to learn node representations preserving local and global network structures for influence maximization, particularly addressing the context of interconnected social networks. Zhang et al. [19] combined convolutional neural networks (CNN) with GNNs in their CGNN algorithm, which simplifies feature matrices and focuses on first- and second-order neighbors to label nodes via the SIR model [59] and optimize a loss function for precise identification of critical nodes. Xiong et al. [16] introduced the AGNN algorithm, which combines autoencoders with GNNs to create topological feature embeddings using GCNs and optimizes ranking predictions via listMLE. Ahmad et al. [60] proposed frameworks (LCNN) that merge CNNs with local node representations and multi-scale metrics. Yu et al. [15] developed RCNN, a method that integrates GCNs with CNN-based adjacency features for efficient critical node detection. Munikoti et al. [61] proposed ILGR, an inductive Graph Neural Network framework that leverages local sub-graph embeddings and a ranking loss to efficiently identify critical nodes and links in large complex networks based on graph robustness metrics. Huang et al. [62] introduced HIVEN, a GNN-based framework designed for heterogeneous information networks (HINs).
These node-ranking methods face several limitations: they rely on fully observable target network topology, overfit to specific structural patterns—hindering cross-network generalization—and often use learning objectives misaligned with the final ranking task. In contrast, ICAN learns robust, low-dimensional embeddings without relying on the topology of target networks, enabling strong generalization to unseen graphs.
Causal representation learning methods: Causal relationships have the ability to capture the fundamental mechanism of data generation and are stable across various contexts [63, 22]. Learning causal representations is significant for enhancing the robustness of predictive models. Zheng et al. [31] introduced NOTEARS, a novel method that reformulates the DAG discovery problem as a continuous optimization task incorporating acyclicity constraints, enabling solution via numerical techniques. Subsequently, Ng et al. [32] expanded upon NOTEARS within a graph autoencoder framework, leveraging multilayer perceptrons (MLPs) to capture and model nonlinear structural equations, thereby advancing the capability to handle complex, non-linear relationships in causal inference. Yang et al. [64] integrated a deep autoencoder and a causal structure learning model to learn causal representations using data. Yu et al. [65] proposed DAG-GNN, a deep generative model that uses a graph neural network-based variational autoencoder to learn DAG structures from data, effectively generalizing linear structural equation models to capture nonlinear relationships and handle diverse variable types.
VI Conclusions
@@
This paper proposes an Influence-aware Causal Autoencoder Network, named ICAN, for node importance ranking in complex networks. The proposed method effectively captures the network-invariant causal relationships between node features and influence scores. This design allows ICAN to learn robust, low-dimensional node embeddings solely from synthetic networks, which can then be applied to node-ranking tasks on various real-world networks. We conduct comprehensive experiments by training ICAN on five types of synthetic networks and evaluating it on diverse benchmark networks, comparing it performance with eleven representative node-ranking methods. The results confirm that ICAN achieves state-of-the-art performance, excelling in both ranking accuracy and cross-network generalization. The results of the ablation studies confirm that both the influence-aware causal representation learning mechanism and the feature–task co-optimization strategy—based on the jointly defined causal ranking loss and causal reconstruction loss—play crucial roles in ensuring that the learned representations generalize effectively across diverse target graphs and are better suited for downstream ranking tasks, thereby substantially enhancing the model’s overall performance. Future work will focus on establishing a theoretical framework for selecting or generating suitable training networks for a given target network, as well as investigating the conditional generalization principle, for the node importance ranking task on one given network.
VII AI-Generated Content Acknowledgment
We did not use any artificial intelligence (AI)-generated content (e.g., text, figures, images, or code) in this article.

### v2

@@ 5.4 Implementation details
The learning rate is $\mu=0.001$ and the maximum number of iterations $T=10$. In addition, we set the recovery probability $\delta=1$ and the infection probability $\gamma=1.5\times\gamma_{c}$ to ensure effective spreading within the network. The infection threshold $\gamma_{c}$ is derived from mean-field theory [moreno2002epidemic] as
@@
For other comparison methods, all the parameters are set to the default. For model evaluation, Kendall’s $\tau$ coefficient on the test networks is adopted as the primary evaluation metric. To reduce the impact of randomness on experimental results, both ICAN and the comparative methods are run 5 times each, with average values recorded. All experiments are conducted on a computer running Windows 10, equipped with an Intel(R) i5-10400 CPU at 2.90 GHz and 16 GB of memory.
5.5 Results on various real-world networks (RQ1)
@@
5.5.1 Training on BA-generated network and testing on various real-world networks
In the first experiment, we conduct experiments by training on the BA generated network and testing on different real-world networks to validate the model’s effectiveness. Table II presents the predicted Kendall’s correlation coefficients across different datasets. The experimental results comprehensively evaluate the proposed ICAN model against several baseline methods for node ranking across six real-world networks. As we can see, ICAN demonstrates superior and consistent performance. Notably, ICAN achieves the highest Kendall’s coefficient on every individual dataset. This culminates in the best overall average Kendall’s coefficient of 0.7707, securing the top rank among all compared methods. The fact that ICAN, trained on a BA synthetic network, generalizes effectively to diverse real-world networks highlights its robust cross-network predictive capability. This can be attributed to the ICAN learning network-invariant node embeddings that transfer reliably across different networks.
@@ 5.5.1 Training on BA-generated network and testing on various real-world networks
5.5.2 Training on ER-generated network and Testing on various real-world networks
@@
Nevertheless, ICAN exhibits relatively lower performance on certain networks, such as USAir and Email-dnc. This performance gap may be largely attributed to the distributional discrepancy between ER training network and the target real-world networks, particularly in terms of degree heterogeneity and community structure. The ER network is characterized by a nearly uniform degree distribution, while many real-world networks display heavy-tailed or scale-free properties, resulting in markedly different local connectivity patterns. Consequently, the model—though effective in capturing causal dependencies among node features—may face challenges in fully adapting to unseen structural regimes. This also motivates our future work to investigate strategies for selecting or generating training networks that better align with the topological properties of specific application domains.
@@ 5.6 Ablation study (RQ2)
5.6.1 Ablation study for the influence-aware causal node embedding mechanism
@@
As shown in the equations of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, both are associated with $\bm{W}$, through which the node embeddings are learned to be causally related to the node influence scores. This indicates that the causal reconstruction loss and the causal ranking loss jointly contribute to the process of influence-aware causal representation learning. To emphasize the advantages of this influence-aware causal mechanism in ICAN, we conduct an ablation study by setting $\lambda_{1}=0$ and fixing $\bm{W}$ as an identity matrix. In this setup, the Markov blanket of the node influence score variable in the causal reconstruction loss is replaced by the low-dimensional node embeddings directly obtained from a traditional autoencoder. This variant can be regarded as a classical autoencoder applied to the ranking task and is denoted as ICAN_w/o_Cau. The comparison results between ICAN and ICAN_w/o_Cau across different networks are shown in Fig. 2(a). A noticeable performance decline of ICAN_w/o_Cau is observed in most node ranking tasks, highlighting the critical importance of the causal mechanism integrated into the autoencoder model.
5.6.2 Ablation study for the feature–task co-optimization mechanism
@@
In this study, we integrate the causal ranking loss and causal reconstruction loss within a unified framework to enable causal feature extraction and ranking tasks. To evaluate the effectiveness of the proposed end-to-end feature–task co-optimization mechanism, we conduct an ablation study by transforming the integrated model into a two-stage implementation, in which the feature extraction and ranking processes are decoupled, thereby removing the synergistic optimization effect. This variant is referred to as ICAN_w/o_Syner. In the first stage, the causal reconstruction loss and a regularization term are employed as the loss function to optimize the causal node embedding module, yielding low-dimensional node representations that comprehensively encapsulate the topological properties of the network. During the second stage, the learned node embeddings are fed into the ranking prediction module, which is optimized using the causal ranking loss along with a regularization term to generate the predicted node influence score. These predictions are then compared against the node influence score derived from the SIR model, with Kendall’s $\tau$ coefficient used as the evaluation metric. The two-stage model is trained on the BA synthetic network and evaluated on six real-world networks. The comparison results between ICAN and ICAN_w/o_Syner on different networks are presented in Fig. 2(b). Performance decline in ICAN_w/o_Syner is observed in most of the node ranking tasks. These results demonstrate that ICAN facilitates mutual reinforcement between causal representation learning and ranking prediction, resulting in more accurate ranking outcomes.
5.6.3 Ablation study for the causal ranking loss
@@
To validate the efficacy of the proposed the causal ranking loss, we conduct ablation experiments, replacing the CausalListMLE with the ListMLE loss for the node importance ranking, defined as ICAN_w/o_CausalRank, and additionally replacing the CausalListMLE with a MSE loss for the node importance score regression, denoted as ICAN_w/o_Rank. As illustrated in Fig. 2(c), the ICAN model consistently and significantly outperforms the ICAN_w/o_CausalRank and ICAN_w/o_Rank variants on all datasets. The superiority of the CausalListMLE loss stems from its inherent ability to capture ordinal relationships within ranked lists, which is more aligned with the nature of the node importance ranking objective. Overall, these findings conclusively demonstrate that integrating the causal ranking loss function is indispensable for achieving high-ranking accuracy, thereby affirming the design rationale behind our task-aware mechanism.
Remark: Our work follows a similar paradigm to the work of AGNN [xiong2024vital] and STGCN [huang2025sparse]. Both frameworks achieve superior node importance ranking by integrating graph neural networks with autoencoder-based representation learning and transfer learning from synthetic to real-world networks. The novelty of our model compared to these two is that we introduce an influence-aware causal node embedding mechanism to ensure that the learned embeddings are causally related to node influence. We further propose a unified objective function that integrates a causal reconstruction loss and a causal ranking loss with a regularization term. This framework synergistically optimizes both feature representation and ranking, which is validated by ablation studies to enhance ranking performance and generalization capability.
5.7 Experiments with various training networks (RQ3)
@@
5.8 Parameter sensitivity analysis
@@
6 Discussion
In this section, we discuss the implications of the proposed model from both network science and engineering perspectives. First, we examine the degree distribution of the most influential nodes identified by the model, providing insights into the structural characteristics that underlie node importance. We then explore the model’s behavior and applicability in social networks with node attributes, highlighting its potential in more complex, attribute-rich environments.
6.1 Degree distribution of Top 10% Influential Nodes
In this subsection, we analyze the degree distribution of nodes identified as important according to the predicted ranking. Figs. 46 compare the degree distribution of the three networks and the corresponding influence scores of the top 10% nodes identified by ICAN, DC, and CGNN. Each subfigure combines a blue histogram representing the network’s degree distribution (left $y$-axis, relative frequency) with red scatter points indicating the influence scores of individual nodes (right $y$-axis, node influence score).
A clear methodological contrast emerges. As expected for the degree–centrality-based approach, the results for DC in Figs. 4(b)–6(b) exhibit a strong positive correlation: nodes with the highest influence scores coincide with those of the largest degrees, concentrating red points at the high-degree end of the spectrum. In contrast, ICAN and CGNN reveal markedly different patterns. High-scoring nodes appear across a broader degree range, with a noticeable concentration in the moderate-degree region. This suggests that these methods attribute high influence to nodes that are not necessarily the network’s primary hubs.
This divergence highlights a key insight: ICAN captures influence patterns that extend beyond simple degree centrality. In network dismantling scenarios, where intervention costs vary across nodes (for example, high-degree nodes may entail greater operational or socio-economic expenses), ICAN’s emphasis on structurally balanced, propagation-relevant nodes results in a prioritized degree distribution conducive to lower-cost, targeted interventions. These findings point to practical strategies for resource-constrained mitigation, including the identification of critical components for the efficient dismantling of non-cooperative networks.
Figure 4: Degree distribution of the detected top $10\%$ important nodes in Jazz network.
Figure 5: Degree distribution of the detected top $10\%$ important nodes in USAir network.
Figure 6: Degree distribution of the detected top $10\%$ important nodes in Email-univ network.
6.2 Case study for social network with attributes
In this case study, we show that ICAN can be applied to a social network with attribute information. The feature matrix can be constructed from node attributes. Cora [cabanes2012cora] is a citation network where nodes represent scientific publications and edges represent directed citation links between them. Each node feature in the Cora dataset is a 1,433-dimensional binary bag-of-words vector representing the presence or absence of specific terms in the paper’s content. Our model is trained on BA synthetic networks (using Node2Vec for feature generation) and evaluated on the attributed Cora real-world network. The experimental results, as shown in Fig. 7, demonstrate that our method achieves superior ranking performance compared to all baseline approaches. The results demonstrate that ICAN can be applied to social networks with attribute information, indicating its potential for broader applications in social network analysis. However, in such attribute-rich networks, inconsistencies may arise between topological structures and node attributes [ma2025node]. Therefore, future work will focus on developing more effective embedding strategies that can integrate both sources of information in a coherent manner.
Figure 7: Kendall’s $\tau$ of different methods on Cora dataset.
7 Conclusions
@@
Quantifying node importance is a challenging task, particularly in non-cooperative scenarios where the target network structure is not accessible in advance due to privacy or security constraints. This paper proposes a node importance ranking method for complex networks based on an influence-aware causal autoencoder model. The proposed method effectively captures the network-invariant causal relationships between node embeddings and influence scores. This design allows ICAN to learn robust, low-dimensional node representations solely from synthetic networks, which can then be applied to node-ranking tasks on various real-world networks. We conduct comprehensive experiments by training ICAN on different types of synthetic networks and evaluating it on diverse benchmark networks, comparing its performance with eleven representative node-ranking methods. The results confirm that ICAN achieves state-of-the-art performance, excelling in both ranking accuracy and cross-network generalization. The results of the ablation studies confirm that both the influence-aware causal node embedding mechanism and the feature–task co-optimization strategy—based on the jointly defined causal ranking loss and causal reconstruction loss—play crucial roles in ensuring that the learned representations generalize effectively across diverse target graphs and are better suited for downstream ranking tasks, thereby substantially enhancing the model’s overall performance. Future work will focus on establishing a theoretical framework for selecting or generating suitable training networks for a given network engineering task. In addition, subsequent research will undertake a deeper investigation of ICAN in practical network engineering contexts, such as developing cost-aware strategies for network dismantling and designing embedding methods that effectively integrate topological and semantic information in social networks.